OpenStudy (anonymous):

Create two sets of three triangle lengths: one set that does create a right triangle and another that does not. Show and explain why the first set does create a right triangle and why the second set does not. Can someone please walk me through this? :/

OpenStudy (anonymous):

They're talking about the side lengths, right? So you could have like {3,4,5} are the side lengths of a right triangle because \(3^2+4^2=5^2\), but {6,9,10} are not the side lengths of a right triangle because \(6^2+9^2\ne10^2\)

OpenStudy (anonymous):

Remember, for any right triangle, \(a^2+b^2=c^2\)

OpenStudy (anonymous):

ohhh okay thanks so much(:(: !